OpenStudy (anonymous):

Ok....you have a barrel of beer and you need to measure out just one gallon. How do you do this if you only have a three-gallon container and a five-gallon container?

OpenStudy (lncognlto):

That's easy. Fill the 3 gal, then fill the 5 gal with the 3 gal. Then refill the 3 gal, fill up the 5 gal, and you have 1 gal left in the 3.

OpenStudy (amistre64):

drink half of it, pour half of it into one of the other container, drink half of whats left, pour half of that into the container ... repeat :)

OpenStudy (amistre64):

\[\sum_1^{\infty}\frac{(-1)^{n+1}}{2^n}=\frac13\]